The Carbondale Hospital is considering the purchase of a new ambulance. The decision will rest partly on the anticipated mileage to be driven next year. The miles driven during the past 5 years are as follows:

Year12345
Mileage3,1004,0503.4003,8003,700

a) Using a 2-year moving average, the forecast for year 6  = _______ miles (round your response to one decimal place).


b) If a 2-year moving average is used to make the forecast, the MAD based on this = _______  miles 


c) The forecast for year 6 using a weighted 2-year moving average with weights of 0.35 and 0.65 (the weight of 0.65 is for the most recent period) =_______  (round your response to the nearest whole number). 

The MAD for the forecast developed using a weighted 2-year moving average with weights of 0.35 and 0.65= _______ miles


d) Using exponential smoothing with α = 0.40 and the forecast for year 1 being 3.100, the forecast for year 6 = _______  miles 
Answer #1

a) using a 2 year moving average method

F(6) = {Actual(4) + Actual(5)}/2

= (3800+3700)/2

= 3750

b) F(3) = (3100+4050)/2

= 3575

F(4) = (4050+3400)/2

= 3725

F(5) = (3400+3800)/2

= 3600

MAD(Mean Absolute Deviation) = Sum of absolute deviations/Number of periods

= (175 + 75 + 100)/3

= 116.7

c) using the weighted average method

F(3) = 0.65*Actual(2) + 0.35*Actual(1)

= 0.65*4050 + 0.35*3100

= 3718

F(4) = 0.65*3400 + 0.35*4050

= 3628

F(5) = 0.65*3800 + 0.35*3400

= 3660

F(6) = 0.65*3700 + 0.35*3800

= 3735

MAD(Mean Absolute Deviation) = (318 + 172 + 140)/3

= 210

d) F(2) = 0.4*Actual(1) + (1-0.4)*F(1)

= 0.4*3100 + 0.6*3100

= 3100

F(3) = 0.4*4050 + 0.6*3100

= 3480

F(4) = 0.4*3400 + 0.6*3480

= 3448

F(5) = 0.4*3800 + 0.6*3448

= 3589

F(6) = 0.4*3700 + 0.6*3589

= 3633